zarkam21:

Help with sqrt y = 8sqrt2

Hero:

Is it \(\sqrt{y} = 8\sqrt{2}\)

Hero:

@zarkam21

Hero:

You can begin by squaring both sides.

zarkam21:

So we are left with y = 8sqrt2^2

Hero:

If you square both sides you get \((\sqrt{y})^2 = (8\sqrt{2})^2\) Then you have to apply the property \((ab)^2 = a^2b^2\) on the RHS to get \(y = 8^2(\sqrt{2})^2 \)

zarkam21:

Y=128

zarkam21:

8^2=64 *2

Hero:

y does equal 128 but \(8^2 \ne 64*2\)

SemiDefinite:

Squaring stuff often involves throwing away information. So: \(\sqrt{y} = 8 \sqrt2 = \sqrt{64} \sqrt{2} = \sqrt{64 \times2} = \sqrt{128}\)